OpenStudy (anonymous):

How do I solve this using substitution? x squared + y squared = 13 2x - y = 4

jimthompson5910 (jim_thompson5910):

first solve the second equation for y 2x - y = 4 -y = 4-2x y = -4 + 2x y = 2x - 4 Then plug this into the first equation x^2 + y^2 = 13 x^2 + (2x-4)^2 = 13 x^2 + 4x^2 - 16x + 16 = 13 I'll let you finish

OpenStudy (anonymous):

Would x = 3\11?

OpenStudy (anonymous):

Or never mind...that's not right

OpenStudy (anonymous):

x = 3 and 1\5

jimthompson5910 (jim_thompson5910):

yes, x = 3 or x = 1/5 are your two solutions for x use them to find the corresponding values of y

OpenStudy (anonymous):

y = -18\5 and 2

jimthompson5910 (jim_thompson5910):

so what are the two ordered pairs

OpenStudy (anonymous):

(3,2) and (1\5, -18\5)

jimthompson5910 (jim_thompson5910):

very good again remember the slashes go the opposite way, but you got it pretty much